40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 745     Entry time: Sun Jul 27 23:06:17 2008
Author: rana 
Type: Update 
Category: PSL 
Subject: PMC, MZ, MC-MMT, etc. 
With the new PMC now in I aligned the MZ to the new beam (there is sadly no steering
between the PMC and the MZ).

I also removed the pickoff that we had put before the MZ just in case we wanted to
move the FSS pickoff to there - its been 2 years now so I guess its not going to happen.


The new PMC's cavity axis seems to be a few hundred microns higher than the old one. So I
tried to move the MZ EOMs to compensate but ended up also steering all of the MZ's mirrors
to get the contrast good, the beam onto the ISS PDs, centered (sort of) onto the MMT lenses
and onto the periscope.

Along the way I also removed some of the vestigial squeezer stuff around the power control
PBS. The output of the PBS now goes directly into the high power dump with no steering. This
eliminated around a dozen clamps, bases, etc. and a couple of mirrors.


The MC is locked on the low power beam we have running through everything. I restored the
PSL launch beam just using the MC-WFS and it locked on a TEM00. So now we know that we
really don't need the PSL quads for this as long as the MC1 angle is stable.

The good news is that the PMC PZT voltage is now flat: the problem must have really been with
the PZT
and not the cabling or notch box like I had wondered about.

Todo:
-----
1) Continue mode matching into the PMC. Its transmission now is around the same as the
   old one.

2) Put a UHV foil covered lead brick onto the PMC to quiet it down.

3) Characterize the PMC loop and retune the body notch for the new body.

4) Tweak the MZ alignment to minimize the RFAM. We can use StochMon to do this as
   long as we have the MC WFS turned off or we can put in a flipper to take the
   beam before the MC and send it to the StochMon RFPD.

5) Re-align onto the ISS.

6) Install irises around the periscope for the beam. The old iris there is way off.

7) Fix PSL ANG and center both POS and ANG.
ELOG V3.1.3-